• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ulmer

    Log entry time 16:09:47 on June11,2002

    Entry number 85895

    keyword=shift summary


    08:38 Start Ela_H0# kinematics (ep scans for angle offsets)
    First run is 12C(e,e') elastic - actually 4.4 MeV state is the "one"
    14:40 Nearly set up for final ep scan point when magnets quench due
    to low coolant flow from ESR. Scott Spiegel was called in; He levels
    are coming back to normal. However, it was deemed too big an
    investment for just one last point, given that we are otherwise ready
    for the beam energy change (from 3 to 5 pass). So that's it.
    Also, now we see a problem with power supply fault on Right arm
    dipole. Scott will investigate while we're in controlled access anyway
    for the OTR problem.

    We got (see runplan for definitions):

    Epscans: Ela_H03, Ela_H04, Ela_H05 and Ela_H06

    12C(e,e') - 4.4 MeV state: Ela_H07